# Reasoning or Memorization? Unreliable Results of Reinforcement Learning Due to Data Contamination

### v1

@@ 4.1 Spurious Rewards on MATH-500
Following the work of Shao et al. (2025), we replicate the performance of Qwen2.5-Math-7B and Llama3.1-8B-Instruct on the MATH-500 benchmark under various reward signal configurations, using the same experimental setup. The training accuracy curves of MATH-500 are illustrated in Figure 3. Interestingly, the results demonstrate that while random rewards and mv-incorrect rewards noticeably boost accuracy for Qwen2.5-Math-7B, they have little or even adverse impacts on the performance of Llama3.1-8B-Instruct. Additionally, we apply the same RLVR procedure to Qwen2.5-Math-7B-Instruct and discover that the resulting gains are marginal when compared with those of Qwen2.5-Math-7B, indicating that the two Qwen variants exhibit differential sensitivity to RLVR.
Considering the base and instruct variants of Qwen are trained under different paradigms: the former is pre-trained as a general language model without exposure to any dialogue-specific templates, whereas the latter undergoes an additional instruction-tuning stage on data wrapped in a fixed dialogue template. This mismatch creates a training–testing gap for the base model at the start point of RLVR, and its initial accuracy is therefore likely underestimated. Consequently, to obtain a fair estimate of each model’s starting point, we next measure performance under four decoding configurations (w/ and w/o templates, with details shown in Table 1), and the corresponding results are summarized in Figure 4. Surprisingly, we discover that applying the official chat template substantially degrades performance for the Qwen base model: both Qwen-2.5-7B and Qwen-2.5-Math-7B suffer pronounced drops once the template is enabled, whether we use greedy or avg@16 sampling.
Building on this observation, we report two additional reference baselines in Figure 3: (i) a red dashed line, denoted Greedy (w/o Template), corresponds to the best performance of the initial model, and (ii) a grey dashed line, denoted pass@16 and adopted from Yue et al. (2025), serves as a plausible upper bound on the models’ initial performance. Viewed against these baselines, the seeming “RL gains” of Qwen-Math-7B largely reflect adaptation to the template format and merely converge to the Greedy (w/o Template) baseline, indicative of memory recall rather than genuine mathematical generalization. However, spurious rewards, e.g., random and mv-incorrect, still boost the accuracy of Qwen base and maintain the performance of Qwen instruct, while degrading Llama-Instruct eventually. This is the point we need to further explore in the following sections.
@@
Model Dataset Size 80%-Problem 60%-Problem 40%-Problem RougeL EM RougeL EM RougeL EM Qwen2.5-Math-7B MATH-500 500 81.25 65.80 78.06 54.60 69.01 39.20 AMC 83 77.38 55.42 70.25 42.17 75.17 36.14 AIME2024 30 74.04 56.67 55.31 20.00 57.72 16.67 AIME2025 30 54.71 16.67 34.88 0.00 27.43 0.00 MinervaMath 272 36.08 2.94 31.22 0.37 29.35 0.00 LiveMathBench 100 42.76 5.00 32.78 0.00 29.97 0.00 Qwen2.5-7B MATH-500 500 66.42 40.20 60.98 21.20 50.36 8.20 AMC 83 73.24 49.40 64.42 33.73 63.79 28.92 AIME2024 30 59.80 30.00 48.69 13.33 44.65 10.00 AIME2025 30 54.61 10.00 37.59 0.00 30.30 0.00 MinervaMath 272 35.24 2.94 32.35 0.37 27.89 0.00 LiveMathBench 100 41.15 4.00 32.74 0.00 27.95 0.00 Llama3.1-8B MATH-500 500 48.33 17.80 40.55 3.80 32.07 0.60 AMC 83 44.54 4.82 30.62 0.00 27.10 0.00 AIME2024 30 50.50 13.33 30.80 0.00 26.08 0.00 AIME2025 30 47.04 10.00 33.49 0.00 25.20 0.00 MinervaMath 272 36.24 2.21 29.52 0.00 27.11 0.00 LiveMathBench 100 35.55 5.00 31.93 0.00 26.88 0.00
4.2 Analysis of Memorization Capability Results
@@
Considering the Qwen series is trained on massive web-scale corpora, we hypothesize that its divergent RLVR behavior from Llama is because the evaluation set MATH-500 may be inadvertently contaminated in Qwen’s large-scale training data, which is hard to eliminate completely. To verify our hypothesis, we probe memorization on several widely used mathematical-reasoning benchmarks. Concretely, we truncate the original questions at 40%, 60%, and 80% of their lengths and feed these partial questions as prompts into the model, and then evaluate the model’s partial-prompt completion rate by computing ROUGE and EM scores (details in 3.2) between the generated completion and ground-truth continuations. In addition, we evaluate the model’s partial-prompt answer accuracy by checking if the continuation contains the correct answer, across both partial and full question settings.
@@
The detailed results are presented in Table 2, revealing strong signs of data contamination in the Qwen2.5 series models when evaluated on commonly used benchmarks, such as MATH-500, AMC, and AIME2024. For instance, when only the first 60% of the questions are provided, Qwen2.5-Math-7B is able to accurately reconstruct more than half of the remaining problems on MATH-500 with greedy decoding, achieving a completion rate of 54.6%. Even when just 40% proportion of the questions are shown, the model still manages to recover 39.2% of the unseen problems on MATH-500. Similar patterns are observed on AMC and AIME2024. These results indicate that the evaluation benchmarks for Qwen2.5 may suffer from data contamination: Although pre-training on massive web-scale corpora brings strong capacity on mathematical reasoning, e.g., superior performance on recently introduced mathematical tasks like LiveMathBench and AIME2025, those large-scale corpora also include publicly available benchmark problems inevitably, leading to less convincing results of old benchmarks, while removing such instances during large-scale crawling is notoriously difficult.
@@
Meanwhile, we summarize the answer accuracy of the model’s output under different proportions of input questions. The detailed results are shown in Figure 5. We observe that under greedy decoding, even when partial questions are provided, the Qwen2.5 models achieve remarkably high accuracy on MATH-500. For instance, with 80% proportion of the questions, Qwen2.5-Math-7B reaches an accuracy of 63.8% on MATH-500. Even when only 40% proportion of the questions are given, the model still achieves 41.2% accuracy. Because our evaluation matches only the final numeric answer, the model can sometimes “accidentally” output the correct value despite faulty reasoning, which explains the anomalous jump observed for Llama on AIME2025 when only 40% of the questions are provided. Among the several questions that the model solves correctly, the responses contain coherent reasoning chains and even syntactically valid Python code, which, however, is not executed. And several examples are shown in Figure 8 to 12. The emergence of such structured solutions (also observed in Shao et al. (2025)) indicates that the training corpora may have included publicly available resources where benchmark problems are accompanied by detailed solutions. Therefore, the sudden performance surge under random rewards or even mv-incorrect rewards, observed in Figure 3, arises because such rewards inadvertently trigger Qwen to retrieve its memorized answers from the pre-training corpus owing to the clipping bias of GRPO (Shao et al., 2025), rather than stimulating Qwen’s existing distribution of reasoning patterns as explained in Shao et al. (2025).
@@ 4.3 Spurious Rewards on RandomCalculation
RandomCalculation Dataset Construction.
@@
To obtain an uncontaminated evaluation benchmark, we employ Algorithm 1 to construct a suite of mathematically challenging yet verifiable computation datasets. These datasets are composed of expressions built from basic numerical elements, including integers from 0 to 100, as well as fractions, squares, and cubes derived from them. Using these components, we randomly generate mathematical expressions involving between 1 and 20 steps, utilizing the four fundamental arithmetic operations: addition, subtraction, multiplication, and division. To construct the final datasets, we append a standardized problem prefix to each generated expression, resulting in 20 sub-datasets, each containing 1,000 unique problems. We refer to this suite of datasets as RandomCalculation. Examples from the datasets can be found in Figure 2.
Qwen2.5’s initial performance on RandomCalculation.
@@
We first test the zero-shot performance of the Qwen2.5 series models on the RandomCalculation dataset, with the results shown in Figure 6. We find that when using chat templates, the models’ mathematical reasoning ability tends to degrade as the number of computation steps increases, leaving ample room for improvement in multi-step calculation problems. When chat templates are removed, the reasoning performance peaks on problems involving three computation steps, and then gradually declines.
RLVR results on RandomCalculation.
@@
We conduct RLVR training on Qwen2.5-Math-7B using the RandomCalculation datasets, aiming to guide the model toward producing answers that are as close as possible to the correct ones. Experiments are set on two datasets comprising 5-step and 10-step calculation problems, respectively. Each dataset contains 1,000 problems, with 700 used for training and the remaining 300 reserved for validation. The training progress is shown in Figure 7, showing that under correct rewards, the model’s performance steadily improves throughout training. With random or incorrect rewards, training becomes unstable and inconsistent. Under inverted rewards, the model collapses rapidly. These findings suggest that for problems not leaked during pretraining, only correct reward signals can effectively guide the model toward improved reasoning performance. For comparison, we also evaluate Llama3.1-8B-Instruct and observe consistent findings: Only accurate reward signals consistently yield stable performance gains.
Qwen’s stronger mathematical capabilities.
@@
Considering obtaining a reward of 1 on a RandomCalculation instance is virtually impossible, we report Max@16, the highest score among 16 samples, instead of Pass@16 in Figure 7. We observe that on our RandomCalculation datasets, Qwen2.5-Math-7B can surpass the Max@16 threshold when provided with correct reward signals. This finding indicates that reward-aligned RLVR effectively transfers the model’s high-accuracy single-step arithmetic skills (see template-enabled accuracy in Figure 6) to more complex multi-step calculations, as illustrated by one reasoning trace in Figure 13. In contrast, under the incorrect and random reward configurations, the Qwen model either maintains its base performance or exhibits only marginal and unstable improvements. This further highlights the critical role of correct and well-aligned reward signals in enhancing model performance on uncontaminated datasets. Notably, Llama3.1-8B-Instruct fails to surpass the Max@16 threshold even when trained with correct reward signals, and its accuracy falls below the greedy-decoding baseline when exposed to spurious signals. This discrepancy indirectly suggests that Qwen2.5 exhibits stronger mathematical capabilities compared to Llama3.1; however, such inherent strength is not the root cause of its performance boosting under spurious reward on contaminated datasets.
5 Discussion
Due to limited computational resources, our experiments were restricted to a subset of commonly used Qwen2.5 series models. The Qwen3 series will be included in our future experimental evaluation. As the successor to Qwen2.5, it is likely that Qwen3 still suffers from similar issues. Many reinforcement learning algorithms have recently been proposed to enhance mathematical reasoning capabilities. Given the rapid pace of development, it is infeasible to conduct a comprehensive evaluation of all these methods in the short term. In future work, our efforts will focus on systematically evaluating a more diverse set of benchmarks, reinforcement learning methods, and model families. In parallel, we will carry out an in-depth theoretical investigation of this unexpected behavior, with the goal of shedding light on its root causes and advancing our understanding of the underlying mechanisms.
@@ 6 Conclusion
Recent studies have shown that the mathematical reasoning performance of Qwen2.5 models can be substantially improved with minimal supervision, ranging from few-shot prompts to single examples or even without additional data. Some studies further suggest that the mathematical reasoning abilities of the Qwen2.5 model can be improved with reward signals that are random or incorrect. While the effectiveness of these methods is striking, their gains appear largely limited to the Qwen2.5 series. This distinctive phenomenon motivated our study, which reveals that the large-scale pretraining corpus of the Qwen2.5 series models may be contaminated by data that overlaps with several widely used mathematical benchmarks, including MATH-500, AMC, and AIME 2024. We suspect that such data contamination is a key factor for Qwen’s improvement on MATH-500 under those spurious rewards. Moreover, to ensure reliable evaluation of RL-based algorithms, we construct a leakage-free RandomCalculation dataset composed of expressions built from basic numerical elements, and a zero-shot test on this dataset shows Qwen has ample room for improvement in multi-step calculation problems. Leveraging this contamination-free dataset, we run carefully controlled RLVR experiments and find that only accurate reward signals deliver consistent performance improvements, while noisy or reversed rewards yield no measurable gains, underscoring the pivotal importance of reward integrity. We therefore recommend that future research conduct evaluations on clean benchmarks and assess multiple model families to ensure their RL-based methods are robust.
@@ Acknowledgments
Appendix A Appendix

### v2

@@ 4.1 Spurious Rewards on MATH-500
Following the work of Shao et al. (2025), we replicate the performance of Qwen2.5-Math-7B and Llama3.1-8B-Instruct on the MATH-500 benchmark under various reward signal configurations, using the same experimental setup. The accuracy curves of MATH-500 are shown in Fig. 3. Interestingly, the results demonstrate that while random rewards and mv-incorrect rewards noticeably boost accuracy for Qwen2.5-Math-7B, they have little or even adverse impacts on the performance of Llama3.1-8B-Instruct. Additionally, we apply the same RLVR procedure to Qwen2.5-Math-7B-Instruct and discover that the resulting gains are marginal when compared with those of Qwen2.5-Math-7B, indicating that the two Qwen variants exhibit differential sensitivity to RLVR under spurious rewards.
Considering the base and instruct variants of Qwen are trained under different paradigms: the former is pre-trained as a general language model without exposure to any dialogue-specific templates, whereas the latter undergoes an additional instruction-tuning stage on data wrapped in a fixed dialogue template. This mismatch creates a training–testing gap for the base model at the start point of RLVR, and its initial accuracy is therefore likely underestimated. Consequently, to obtain a fair estimate of each model’s starting point, we next measure performance under four decoding configurations as shown in Tab. 1. The corresponding results are summarized in Fig. 4. Surprisingly, we discover that applying the official chat template substantially degrades performance for the Qwen base model: both Qwen-2.5-7B and Qwen-2.5-Math-7B suffer pronounced drops once the template is enabled.
@@
Building on this observation, we report two additional initial accuracy for reference in Fig. 3: (i) Greedy (w/o Template) corresponds to the best performance of the initial model, and (ii) pass@16, adopted from Yue et al. (2025), serves as a plausible performance upper bound for the initial model. Viewed against these baselines, the seeming ‘RL gains’ of Qwen-Math-7B largely reflect adaptation to the template format and merely converge to the Greedy (w/o Template) baseline, indicative of memory recall rather than genuine mathematical generalization. However, spurious rewards, e.g., random and mv-incorrect, still boost the accuracy of Qwen base and maintain the performance of Qwen instruct, while degrading Llama eventually. This is the point we need to further explore in the following sections.
[…]
4.2 Analysis of Memorization Capability
@@
Considering the Qwen series is trained on massive web-scale corpora, we hypothesize that its divergent RLVR behavior from Llama is because the evaluation set MATH-500 may be inadvertently contaminated in Qwen’s large-scale training data, which is hard to eliminate completely. To verify our hypothesis, we probe memorization on several widely used mathematical-reasoning benchmarks. Concretely, we truncate the original questions at 40%, 60%, and 80% of their lengths and feed these partial questions as prompts into the model, and then evaluate the model’s partial-prompt completion rate by computing ROUGE and EM scores between the generated completion and ground-truth continuations. In addition, we evaluate the model’s partial-prompt answer accuracy by checking if the continuation contains the correct answer, across both partial and full question settings.
@@
The detailed results are presented in Tab. 2, revealing strong signs of data contamination in the Qwen2.5 series models when evaluated on commonly used benchmarks, such as MATH-500, AMC, and AIME2024. For instance, when only the first 60% of the questions are provided, Qwen2.5-Math-7B is able to accurately reconstruct more than half of the remaining problems on MATH-500. Even when just 40% proportion of the questions are shown, the model still manages to recover 39.2% of the problems on MATH-500. Similar patterns are observed on AMC and AIME2024. These results indicate that the evaluation benchmarks for Qwen2.5 may suffer from data contamination. Although pre-training on massive web-scale corpora brings strong capacity on mathematical reasoning, e.g., superior performance on recently introduced mathematical tasks like LiveMathBench and AIME2025, those large-scale corpora also include publicly available benchmark problems inevitably, leading to less convincing results of old benchmarks, while removing such instances during large-scale crawling is notoriously difficult.
@@
Meanwhile, we summarize the answer accuracy of the model under different ratios of prefix in Fig. 5. The Qwen2.5 models achieve remarkably high accuracy on MATH-500 even with partial questions. For instance, with 80% proportion of the questions, Qwen2.5-Math-7B reaches an accuracy of 63.8% on MATH-500. Even with only 40% proportion of the questions, the model still achieves an accuracy of 41.2%. Because our evaluation matches only the final numeric answer, the model can sometimes output the correct value by accident, which explains the anomalous accuracy of Llama on AIME2025 (solving exactly one problem with a faulty reasoning process). Besides, we also inspect several questions that Qwen solves correctly, as shown in Fig. 10 to 14 in Appendix. We find that the responses of Qwen contain coherent reasoning chains and even syntactically valid Python code, which, however, is not executed. The emergence of such structured solutions indicates that the training corpora may have included publicly available resources where benchmark problems are accompanied by detailed solutions. Besides, we provide more results for Qwen2.5 and Qwen3 under various sampling configurations in Appendix. The memorization of Qwen3 series is also analyzed in Appendix, showing similar results as Qwen2.5.
@@ 4.3 Spurious Rewards on RandomCalculation
Dataset Construction of RandomCalculation.
@@
To obtain an uncontaminated evaluation benchmark, we employ Algorithm 1 (shown in Appendix) to construct a suite of challenging yet verifiable datasets. These datasets are composed of expressions built from basic numerical elements, including integers from 0 to 100, as well as fractions, squares, and cubes derived from them. Using these components, we randomly generate mathematical expressions that involve between 1 and 20 steps, using the four fundamental arithmetic operations: addition, subtraction, multiplication, and division. To construct the final datasets, we append a standardized problem prefix to each generated expression, resulting in 20 sub-datasets, each containing 1,000 unique problems. We refer to this suite of datasets as RandomCalculation. Examples from the datasets can be found in Fig. 2.
Zero-shot Performance on RandomCalculation.
@@
We first test the zero-shot performance of the Qwen2.5 series on the RandomCalculation dataset in Fig. 6. We find that when using chat templates, the models’ performance degrades gradually as the number of computation steps increases, leaving ample room for improvement in multi-step calculation problems. When chat templates are removed, the reasoning performance peaks on problems of three computation steps, and then gradually declines.
Correct Reward Function for RandomCalculation.
The ground-truth answers to our randomly generated arithmetic problems often contain high-precision decimals. When using the standard RLVR framework, which only provides binary feedback, the model rarely receives positive reinforcement, making training unstable and prone to divergence. To overcome this, we design a continuous reward function that ranges from 0 to 1 and penalizes both absolute and relative errors between the model’s prediction and the reference answer. This richer feedback helps stabilize reinforcement learning. Let be the model output, be the reference answer, and be a small constant for numerical stability. The reward is computed as:
RLVR on RandomCalculation.
@@
We also perform RLVR training on Qwen2.5-Math-7B using the RandomCalculation datasets. Specifically, experiments are conducted on two sub-datasets comprising 5-step and 10-step calculation problems. Each dataset contains 1,000 problems, with 700 used for training and the remaining 300 reserved for validation. As shown in Fig. 7, the performance improves steadily throughout training under correct rewards. However, training becomes unstable and inconsistent with random or incorrect rewards. Under inverted rewards, the performance collapses rapidly. These findings suggest that for problems not leaked during pretraining, only correct reward signals can effectively guide the model toward improved performance. For comparison, we also evaluate Llama3.1-8B-Instruct and observe similar findings.
Qwen v.s. Llama on Clean Benchmark.
@@
Considering obtaining a reward of 1 on a RandomCalculation instance is virtually impossible, we report Max@16, the highest reward among 16 samples of initial model, in Fig. 7. We observe that on RandomCalculation datasets, Qwen2.5-Math-7B can surpass Max@16 when provided with correct reward signals. This finding indicates that reward-aligned RLVR effectively transfers the high-accuracy single-step arithmetic skills (as shown in Fig. 6) to more complex multi-step calculations, as illustrated by one reasoning trace in Fig. 15 of Appendix. In contrast, under the incorrect and random reward configurations, the Qwen model either maintains its base performance or exhibits only marginal and unstable improvements, due to the learning of format as explained in Shao et al. (2025). This improvements gradually disappear when calculation steps increased from 5 to 10. This further highlights the critical role of correct and well-aligned reward signals in enhancing model performance on uncontaminated and difficult datasets. Notably, Llama3.1-8B-Instruct fails to surpass the Max@16 even when trained with correct reward signals, and its accuracy falls below the greedy-decoding baseline when exposed to spurious signals. This discrepancy indirectly suggests that Qwen2.5 exhibits stronger mathematical capabilities compared to Llama3.1. However, such inherent strength is not the root cause of its performance boosting under spurious reward on contaminated datasets.
4.4 More Evidence for Memorization
Here, we provide more detailed analyses of Qwen’s sudden performance gains on MATH-500 under random reward. Let where is a random variable under the setup of random reward. Referring to Appendix B of Shao et al. (2025), the gradient of the clipped policy has the following format:
where is a positive coefficient, .
Memory Retrieval due to Exploitation Bias.
Assume a high-probability token with and , the upper clipping boundary is for , which exceeds the probability ceiling of and therefore is never reached. Consequently, the gradient is non-negative for this token, leading to a net positive gradient bias on the policy model, In general, for high-probability token, we have , due to hold almost surely. So high‐probability tokens continue to be up-weighted without penalty. For MATH‐500, correct answers typically have a high probability due to data contamination in the initial model (results shown in Fig.9 of Appendix), except for the low-probability answer format. Therefore, GRPO with random reward can retrieve these answers after learning format and leads to sharp accuracy jump in Fig 3.
On the other hand, assume another token with pre-update likelihood (which is a typical value for our 10-step RandomCalculation as shown in Fig. 9 of Appendix). The corresponding clipping boundary is for . Gradient update with random reward perturbs around this narrow band, so that in most cases. Consequently, Therefore, no meaningful performance improvement observed in 10-step RandomCalculation with random reward in Fig. 7. Overall, clipped objective introduces systematic exploitation bias for high-probability tokens, whereas mid-probability tokens are less optimized.
Response Similarity Before and After RL.
We further compare the responses of the model before and after RL, with Rouge-L and KL distance111For each generated answer we compute over the full vocabulary, where and denote probability distributions produced by the fine-tuned and base models, respectively. as the similarity score. As shown in Tab. 3 and Fig. 8, the similarity in MATH-500 is substantially higher than in RandomCalculation, further implying that MATH-500 suffers from data contamination. Additionally, spurious rewards achieve even higher Rouge-L than correct reward after RL. Therefore, performance surge under spurious rewards arises because GRPO inadvertently triggers Qwen to retrieve memorized answers, rather than stimulating Qwen’s existing reasoning patterns like codes as explained in Shao et al. (2025). This is due to the exploitation bias of GRPO. However, RL with correct reward can still stimulate the model to find new reasoning paths, as verified with a smaller Rouge-L of response.
| Dataset | Reward Signal | ROUGE-L |
| MATH-500 | Correct | 0.555 |
| Random | 0.601 | |
| Mv-Incorrect | 0.563 | |
| RandomCalculation 5 Steps | Correct | 0.225 |
| Random | 0.247 | |
| Mv-Incorrect | 0.251 | |
| RandomCalculation 10 Steps | Correct | 0.193 |
| Random | 0.251 | |
| Mv-Incorrect | 0.279 |
@@ 5 Conclusion
In this work, we investigate the unexpected performance improvements of Qwen on mathematical reasoning with spurious rewards. Our analysis reveals that these gains were primarily due to data contamination rather than Qwen’s inherent mathematical capabilities. By auditing the MATH-500 dataset and introducing a clean benchmark, we demonstrate that Qwen’s successes with spurious reward were driven by memorization of benchmark problems rather than genuine reasoning skills. Additionally, we show that only correctly aligned rewards lead to consistent performance improvements, while spurious rewards fail to provide meaningful benefits. These findings underscore the importance of using uncontaminated benchmarks in evaluating RL-based methods and call for caution when interpreting results from datasets that may suffer from data leakage. Our work highlights the need for rigorous evaluation protocols in future research to ensure that performance gains reflect true advancements in ability, rather than data contamination.
@@ Acknowledgments
Appendix A Discussion and Limitation
Due to limited computational resources, our experiments were restricted to a subset of commonly used Qwen2.5 and Qwen3 series. Besides, given the rapid development of various RL algorithms, it is infeasible to conduct a comprehensive evaluation of all these methods in the short term. In future work, our efforts will focus on comprehensive evaluation on more diverse benchmarks, reinforcement learning methods, and model families.
Appendix B Details of RandomCalculation Construction
We provide the specific algorithm for the construction of our clean RandomCalculation benchmark here:
Appendix C Quantitative Analysis of Answer-Related Numeric Token Probabilities
Fig. 9 illustrates the change in probabilities of numeric tokens associated with final answers before and after reinforcement learning (RL), evaluated on 100 samples from the MATH-500 and 10-step RandomCalculation datasets using random reward. The MATH-500 results exhibit consistently high probabilities both before and after RL, suggesting that the model retains a strong memory of the answers. In contrast, the results on the clean 10-step RandomCalculation dataset are more dispersed, showing that random reward signals are ineffective.
Appendix D Quantitative Results on Partial-Prompt Answer Accuracy
Tab. 4 presents an expanded comparison of partial-prompt answer accuracy between Qwen-2.5 and Llama-3.1 on the MATH-500 benchmark. To complement these findings, Tab. 5–8 reports analogous results across several additional mathematics datasets, evaluated under four generation configurations: Greedy decoding with and without templates, and Avg@16 decoding with and without templates.
Appendix E Quantitative Results for Qwen3
Qwen3 is the closest open-source model within the Qwen family, and we subject it to the same memorization diagnostics. Table 9 reports its partial-prompt completion rate, whereas Tables 10–13 present the corresponding partial-prompt answer accuracy under all generation settings. The results mirror those of Qwen2.5: despite the model’s increased capacity, it still exhibits pronounced evidence of data contamination, likely attributable to pre-training on large-scale web corpora.
Appendix F Illustrative Examples of Model Output
Figures 10 and 11 present concrete instances of memorization by Qwen-2.5-7B on the AMC and AIME2024 benchmarks, respectively. Figures 12–14 provide analogous examples for Qwen-2.5-Math-7B on MATH-500, AMC, and AIME2024. Figure 15 depicts a representative response produced by the RLVR-fine-tuned Qwen-2.5-Math-7B on the 5-step RandomCalculation benchmark.